# Information theory for data-driven model reduction in physics and biology

### v2

@@ Abstract
Model reduction is the construction of simple yet predictive descriptions of the dynamics of many-body systems in terms of a few relevant variables. A prerequisite to model reduction is the identification of these relevant variables, a task for which no general method exists. Here, we develop a systematic approach based on the information bottleneck to identify the relevant variables, defined as those most predictive of the future. We elucidate analytically the relation between these relevant variables and the eigenfunctions of the transfer operator describing the dynamics. Further, we show that in the limit of high compression, the relevant variables are directly determined by the slowest-decaying eigenfunctions. Our information-based approach indicates when to optimally stop increasing the complexity of the reduced model. Furthermore, it provides a firm foundation to construct interpretable deep learning tools that perform model reduction. We illustrate how these tools work in practice by considering uncurated videos of atmospheric flows from which our algorithms automatically extract the dominant slow collective variables, as well as experimental videos of cyanobacteria colonies in which we discover an emergent synchronization order parameter.
Model reduction |||| Information theory |||| Dynamical systems
@@
The exhaustive description of a biological or physical system is usually impractical due to the sheer volume of information involved. As an example, the air in your office may be described by a $10^{27}$–dimensional state vector containing the positions and momenta of every particle in the room. Yet, for most practical purposes, it can be effectively described using only a small number of quantities such as pressure and temperature. Similar reductions can be achieved for systems ranging from diffusing particles to biochemical molecules and complex networks. In all cases, certain relevant variables can be predicted far into the future even though individual degrees of freedom in the system are effectively unpredictable.
@@
The process by which one goes from the complete description of a system to a simpler one is known as model reduction. Diverse procedures for model reduction exist across the natural sciences. They range from analytical methods, such as adiabatic elimination and multiple-scale analysis [1, 2, 3, 4, 5, 6, 7, 8, 9], to data-driven methods such as independent component analysis [10], dynamic mode decomposition [11, 12, 13], diffusion maps [14], spectral submanifolds [15, 16], and deep encoder-decoder neural networks [17, 18, 19, 20, 21, 22, 23, 24].
The success of these approaches is limited by a fundamental difficulty: before performing any reduction, one has to identify a decomposition of the full system into relevant and irrelevant variables. In the absence of prior knowledge and intuition (e.g. a clear separation of scales), identifying such a decomposition is an open problem [4]. It may not even be clear a priori when to stop increasing the complexity of a simplified model or, conversely, when to stop reducing the amount of information needed to represent the dynamical state of a complex system. In both cases one must first determine the minimal number of relevant variables that are needed. The answer to this question depends in fact on how precisely and how far in the future you wish to forecast. Nonetheless, this answer should be compatible with fundamental constraints on forecasting set by external perturbations and finite measurement accuracy [25, 26].
@@
In order to address the difficulty identified in the previous paragraph, we develop an information-theoretic framework for model reduction. Very much like MP3 compression is about retaining information that matters most to the human ear [27], model reduction is about keeping information that matters most to predict the future [28, 29]. Inspired by this simple insight, we formalize model reduction as a lossy compression problem known as the information bottleneck (IB) [30, 31]. This formal step allows us to give a precise answer to the question of how to identify relevant and irrelevant variables. We show how and under what conditions the standard operator-theoretic formalism of dynamical systems [32, 19], which underlies most methods of model reduction, naturally emerges from optimal compression. Crucially, our framework systematically answers the question of when to stop increasing the complexity of a minimal model. Further, it provides a firm foundation to address a practical problem: the construction of deep learning tools to perform model reduction that are guaranteed to be interpretable. We illustrate our approach on benchmark dynamical systems and demonstrate that it works even on uncurated datasets, such as satellite movies of atmospheric flows downloaded directly from YouTube and biological datasets composed of microscopy videos of cyanobacteria colonies in which we discover an emergent synchronization order parameter.
Figure 1: Interpretable dynamical variables in reduced models via the information bottleneck. (a) The information bottleneck compresses high-dimensional state variables $x_{t}$, into simpler encoding variables $h_{t}$ providing a controllable trade-off between the degree of compression and the predictive power about the system’s future. With deep neural networks, the encoding can be computed directly from data of observed fluid flows (left) or biological datasets, such as fluorescently labeled bacteria colonies (right). In general, the state of the variable $x_{t}$ may comprise time-lagged variables of the intensity field, $x_{t}=\{I_{t},I_{t+\Delta t}\}$ (right). The amount of compression is determined by the “width” of the bottleneck $\beta$ [see (1))]. The resulting compressed, or encoded, variables $h_{t}$ represent collective variables most predictive of the system’s future. (b) Schematic evolution of the encoder $p(h_{t}|x_{t})$ for varying compression strength $\beta$. For low $\beta$ (high compression), the encoder is trivial and forgets everything about the input $x_{t}$. After the first IB transition at $\beta_{1}$, the encoder becomes non-trivial by gaining some dependence on $x_{t}$; some features of the input are able to pass through the bottleneck. At subsequent IB transitions, additional features are learned. (c) The point spectrum of the transfer operator contains several slowly decaying modes (red). We show that the most predictive variables that IB systematically extracts correspond to the slowest eigenfunctions of the transfer operator, associated to eigenvalues $\Lambda_{i}$ with $|\Lambda_{i}|\approx 1$. In fluid flows, the slowest-decaying eigenfunctions typically represent large-scale coherent patterns of the flow field, while faster-decaying eigenfunctions correspond to variations over shorter length scales.
@@ I Model reduction as a compression problem
We present here a method to extract collective variables most predictive of the system’s future evolution directly from data. This data is composed of a time sequence of measured states $x_{1},x_{2},...,x_{T}$. The system state $x_{t}$ could correspond to anything from the position of a single particle to an image of a fluid flow or the fluorescent molecules in a living system (Fig. 1a). The full state can be very high dimensional, with a number of dimensions equal to the number of observed pixels in the case of imaging data. However, the variation of any individual pixel is often of limited interest to us, as noise (either inherent or due to measurement) induces uncertainty about its true value. Individual pixels are, due to this uncertainty, poor predictors of the future state of the system. We can say that they are irrelevant for predicting the future. On the other hand, certain spatially-averaged collective variables may evolve slowly in time, and the future state of the system may be reliably estimated from them.
We seek a way to “encode” each state $x_{t}$ into a simple, lower-dimensional representation $h_{t}$ in a way that isolates these relevant features of the input $x_{t}$. For instance, in Fig. 1a, both the velocity field of a fluid flow and the images of a dynamic cyanobacteria colony (upper row) may be encoded as a point in a 2D space (lower row). The encoding is given by a probabilistic mapping $p(h_{t}|x_{t})$ which can be thought of as a machine which takes a state $x_{t}$ and assigns it to a value $h_{t}$. The fact that this mapping is probabilistic simply means we may have some uncertainty about the true value of $h_{t}$ even given a measurement of the state. Whether or not some encoding is extracting “relevant” features of the state $x_{t}$ is determined by the extent to which we can use it to predict the future. This predictive power can be quantified by the mutual information between the encoding and the future state, $I(H_{t},X_{t+\Delta t})$, which tells us how much the knowledge of $H_{t}$ reduces our uncertainty about the future $X_{t+\Delta t}$ (see Methods). (In our notation, upper-case $X_{t}$ refers to the random variable, while $x_{t}$ refers to a particular value taken by the random variable.) To find a good encoder that extracts relevant features, we might try to find an encoding which maximizes this information. However, an encoder obtained in this way would simply copy the original state, $h_{t}=x_{t}$, since $x_{t}$ represents all the information we have about the system. In order to encourage the encoder to discard irrelevant features, we simultaneously seek an encoder which maximizes compression by minimizing the information about the original state, $I(H_{t},X_{t})$. This prescription for encoding relevant collective variables can be formalized by the information bottleneck (IB) method [30, 28, 31]. The information bottleneck objective function combines both of our stated goals – compression and predictive power – into one mathematical expression:
$\mathcal{L}_{\text{IB}}[p(h_{t}|x_{t})]$ $=I(X_{t},H_{t})-\beta I(X_{t+\Delta t},H_{t}).$ (1)
@@
Crucially, the parameter $\beta$ allows one to tune how much weight to assign to compression versus prediction. For small $\beta$ the compression term dominates and the optimal encoder is trivial, losing all information about the system. For intermediate $\beta$ the compression term does not allow $X_{t}$ to be completely captured by $H_{t}$, so that features of $X_{t}$ must “compete” to pass through to the encoding variable (Fig. 1b). These features are reflected in the form of the encoder
@@
which provides the optimal trade-off between compression and predictability [29]. Our goal is to connect the dynamical properties of the system to the features learned by the encoder.
@@
In any realistic experimental setting, the presence of noise or uncertainty means we cannot predict precisely the future state of a system but instead can only predict a likely distribution of possible future states. Our prediction of the state at $\Delta t$ in the future is then represented mathematically as $p(x_{t+\Delta t}|x_{t})$, the probability of observing state $x_{t+\Delta t}$ given the current state $x_{t}$. This conditional probability distribution completely characterizes the dynamics of the system, and determines how probability distributions evolve in time:
$p(x_{t+\Delta t})=\int p(x_{t+\Delta t}|x_{t})p(x_{t})\text{d}x_{t}.$ (3)
For Markovian, or “memoryless” dynamics, such an evolution can be understood as the action of a (linear) transfer operator $U$ which acts on probability distributions. $U$ can be decomposed into its right and left eigenvectors as
$U=\sum_{n}\ket{\rho_{n}}\text{e}^{\lambda_{n}\Delta t}\bra{\phi_{n}}+U_{\text{ess}}$ (4)
where |$\ket{\rho_{n}}$ are right eigenvectors with eigenvalue $\Lambda_{n}\equiv\text{e}^{\lambda_{n}\Delta t}$ and $\bra{\phi_{n}}$ | are the corresponding left eigenvectors. $\lambda_{n}$ are the eigenvalues of the infinitesimal generator of $U$, known as the Fokker-Planck operator (Fig. 1c). The operator $U_{\text{ess}}$ corresponds to the so-called essential spectrum, and we assume that it can be neglected. This is usually possible when the system is subjected to even a small amount of noise, or when some amount of uncertainty is present in the measurements [33, 34]. The eigenfunctions $\phi_{n}$ in (4) are in some sense “natural” features of the dynamics, as they evolve independently in time.
@@
Our key observation is that the optimal encoder in (2) can be expressed in terms of the eigenvalues $\lambda_{n}$ and left eigenfunctions $\phi_{n}$ of (the generator of) $U$,
@@
where $f_{n}(h_{t})$ are factors that do not depend on $x_{t}$. For an outline of the mathematical steps leading to this see Methods, as well as the SI. These factors effectively determine what the encoder learns about the state $x_{t}$. In general, there may be a large number of non-zero factors $f_{n}$ so that the learned features are difficult to extract. However, things become simple in the limit of small $\beta$, or high compression. When $\beta$ is small the encoder is trivial: $p(h_{t}|x_{t})=p(h_{t})$. In this case the value $h_{t}$ is assigned at random with no regard to the state $x_{t}$ of the system. No feature has been learned, and all factors $f_{n}$ are equal to zero. As $\beta$ is increased, the encoder undergoes a series of transitions at $\beta=\beta_{1}<\beta_{2}<\beta_{3}...$ where new features are allowed to pass through the bottleneck (Fig. 1b) [35, 36, 37, 38]. The first transition happens at a finite value of $\beta_{1}$ when the first most predictive feature is learnt.
@@
This is our main mathematical result, which we derive by considering a perturbative expansion of the IB objective for small $f_{n}$. A proof of Eq. 6 with clearly specified technical assumptions may be found in the Methods and SI.
@@
The above statement shows that in the limit of high compression the encoder’s dependence on $x_{t}$ is given by the first left eigenfunction $\phi_{1}(x_{t})$, which is the slowest-varying function of the state under dynamics given by $U$. Therefore, Eq. 6 makes precise the intuitive statement that slow features are the most relevant for predicting the future. Our analytical result, while applying only to the dominant eigenfunction, is valid for arbitrary non-Gaussian variables. The question of maximally informative features has additionally been explored in the context of animal vision, where one seeks to understand what features of the field of vision are encoded by retinal neurons [39, 40].
@@
We further observe numerically that this picture holds true more generally: also at successive IB transitions, the learned features correspond to successive modes of the transfer operator. This picture is consistent with the exact results known for Gaussian IB, where the encoder learns eigenvectors of a matrix (related to the covariance of the joint $X_{t},X_{t+\Delta t}$ distribution) in a step-wise fashion at each IB transition [37]. Together, this shows that the most informative features extracted by IB, an agnostic information-theoretic approach, correspond to physically-interpretable quantities – namely transfer operator eigenfunctions. As we show later, the insight above can be leveraged to systematically learn these slow variables directly from data with neural networks [41].
II Information decay and the spectrum of the transfer operator

### v3

@@ Abstract
Model reduction is the construction of simple yet predictive descriptions of the dynamics of many-body systems in terms of a few relevant variables. A prerequisite to model reduction is the identification of these variables, a task for which no general method exists. Here, we develop an approach to identify relevant variables, defined as those most predictive of the future, using the so-called information bottleneck. We elucidate analytically the relation between these relevant variables and the eigenfunctions of the transfer operator describing the dynamics. In the limit of high compression, the relevant variables are directly determined by the slowest-decaying eigenfunctions. Our results provide a firm foundation to interpret deep learning tools that automatically identify reduced variables. Combined with equation learning methods this procedure yields the hidden dynamical rules governing the system’s evolution in a data-driven manner. We illustrate how these tools work in diverse settings including model chaotic and quasiperiodic systems in which we also learn the underlying dynamical equations, uncurated satellite recordings of atmospheric fluid flows, and experimental videos of cyanobacteria colonies in which we discover an emergent synchronization order parameter.
Model reduction $|$ Information theory $|$ Dynamical systems
@@
The exhaustive description of a biological or physical system is usually impractical due to the sheer volume of information involved. As an example, the fluids in a steam engine may be described by a $10^{27}$–dimensional state vector containing the positions and momenta of every particle in the engine. Yet, for most practical purposes, its working can be effectively described using only a small number of thermodynamic variables like pressure and temperature. Likewise, some aspects of the behavior of a nematode or a fruit fly, with $385$ and $\sim$140,000 neurons respectively, may be effectively captured by orders of magnitude fewer degrees of freedom [1, 2, 3, 4, 5]. Similar reductions can be achieved for systems ranging from diffusing particles to biochemical molecules and complex networks. In all cases, certain relevant variables can be predicted far into the future even though individual degrees of freedom in the system are effectively unpredictable.
@@
The process by which one goes from the complete description of a system to a simpler one is known as model reduction. Diverse procedures for model reduction exist across the natural sciences. They range from analytical methods, such as adiabatic elimination and multiple-scale analysis [6, 7, 8, 9, 10, 11, 12, 13, 14], to data-driven methods such as dynamic mode decomposition and other linear projection-based methods [15, 16, 17, 18, 19, 20], diffusion maps [21], spectral submanifolds [22, 23], independent component analysis [24], and deep encoder-decoder neural networks [25, 26, 27, 28, 29, 30, 31, 32].
Model reduction consists of two ingredients: a decomposition of the full system into relevant and irrelevant variables, and a self-contained dynamical equation which governs the evolution of the relevant variables. In the absence of prior knowledge and intuition (e.g. a clear separation of scales), identifying a decomposition that produces a minimal set of variables which are maximally descriptive of the full system is an open problem [9]. The answer to this question depends both on the time scale and precision with which one aims to describe the system [33, 34]. In addition, it also depends on one’s ability to identify an adequate dynamical equation for the relevant variables, especially in data-driven settings, where the relevant variables and the dynamical equation are to be learned simultaneously.
To make progress, we develop an information-theoretic framework for model reduction. Very much like MP3 compression is about retaining information that matters most to the human ear [35], model reduction is about keeping information that matters most to predict the future [36, 37]. However, formalizing this intuitive statement requires some care in order to avoid trivial results, that manual approaches usually avoid using physical intuition. By carefully formulating model reduction within this information-theoretic perspective, we show that model reduction (the identification of both relevant variables and their effective dynamics) can be decomposed into a two-step process. First, the problem of variable identification can be reformulated in a way that removes any reference to the effective dynamics, and solved on its own. Second, once the relevant variables are known, an independent dynamical-equation learning step can yield the effective dynamics.
@@
Our key step is to relate the model reduction objective to a lossy compression problem known as the information bottleneck (IB)[38, 39, 36], and to analytically show how and under what conditions the standard operator-theoretic formalism of dynamical systems [40, 27], which underlies most methods of model reduction, naturally emerges from the optimal compression solution. This allows us to give a precise answer to the question of how to identify relevant and irrelevant variables. Crucially, our framework provides a criterion for when to stop increasing the complexity of a minimal model by directly measuring how informative relevant variables are of their own future and of the future of the entire system. Further, it provides a firm foundation to address a practical problem: the construction of deep learning tools to perform model reduction that are guaranteed to be interpretable. This practical realization of our protocol is based on recently-developed deep learning tools that allow scaling our approach to high dimensional systems [41].
The remainder of this paper is organized as follows. Section 1 formalizes model reduction and gives our information-theoretic starting point (see also SI Sections 1-3). Section 2 shows how the optimal information-bottleneck solution identifies the dominant (left) eigenfunctions of the system’s transfer operator (see also SI Section 4), and delineates a procedure to extract the reduced variables. In Sections 3 and 4 we show how information measures capture features of the system’s transfer operator spectrum and how this relates to the question of “when to stop” increasing the complexity of the reduced model. Section 5 (and SI Sections 5-6) shows that our framework can be used to learn relevant variables directly from data using neural networks. We illustrate our approach on several benchmark dynamical systems and demonstrate that it works even on uncurated datasets, e.g. satellite movies of atmospheric flows downloaded from YouTube. SI Section 7 shows applications to chaotic dynamics including the Lorenz-63 system, the Kuramoto-Sivashinsky equation and a hyperchaotic system. In Section 6 we show how our approach, combined with equation learning algorithms, forms a full model reduction pipeline yielding both the reduced variables describing the system and the equations governing its dynamics. Finally, in Section 7 (and SI Section 8) we apply our method to experimental microscopy videos of cyanobacteria colonies, discovering an emergent synchronization order parameter.
@@ I Model reduction as a compression problem
Consider a system whose state is described by a (random) variable $X_{t}$, which might correspond to anything from the position of a single particle to an image of a fluid flow or the fluorescent molecules in a living system (Fig. 1a,b). The full state can be high dimensional, with a number of dimensions equal to the number of particles in a gas, or the number of pixels in an image. It is often difficult to predict the future state $X_{t+\Delta t}$ of such high-dimensional systems due to the enormous computational cost of evolving the system forward in time to a desired accuracy. However, in many cases there exist some functions of the full state $H_{t}=h(X_{t})$ which are low-dimensional and very informative of the future state of the system. These features often correspond to collective variables which evolve slowly in time. While one may guess the appropriate choice of variables $h(X_{t})$ in contexts where all microscopic details of the system are known, we are interested in how to find such variables without such a priori knowledge.
We formulate the task of identifying the reduced variables and their evolution as an optimization problem. A “good” reduced model is one which entails a compressed description of the state, and which may be evolved in time in place of the full system to make predictions. The procedure consists of three steps: first, reduction of the full system by an encoding $h_{\text{enc}}\mathrel{\mathop{\ordinarycolon}}X_{t}\mapsto H_{t}$; second, evolution of the reduced state via an evolution operator $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}H_{t}\mapsto H_{t+\Delta t}$; third, decoding the reduced state via a function $h^{-1}_{\text{dec}}\mathrel{\mathop{\ordinarycolon}}H_{t+\Delta t}\mapsto X_{t}$ to predict the full system’s state. These functions can be found by optimizing
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}$ $\dim h_{\text{enc}}(X_{t})+\beta\,d\big(h^{-1}_{\text{dec}}(\mathcal{S}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big).$ (1)
The first term is a compressive constraint enforcing a “reduction” of the model, with the $\dim(\cdot)$ function measuring e.g. the number of bits used to represent the encoded state, or the sparsity of its representation in a preferred basis. The second term captures the error in our prediction of the full state when using the reduced model, measured by a “distance” $d(\cdot,\cdot)$. The parameter $\beta$ sets the trade-off between dimensionality reduction and predictive power. Note that if the encoding $h_{\text{enc}}$ is non-invertible (as is generally the case), the decoder $h^{-1}_{\text{dec}}$ is defined probabilistically via a conditional distribution $h^{-1}_{\text{dec}}(H_{t+\Delta t})\sim p(\cdot|H_{t+\Delta t})$. This is analogous to a procedure producing a thermodynamic ensemble compatible with the state variables. For instance, the canonical ensemble associates to every macrostate $h_{\text{enc}}=(N,V,T)$ defined by the number of particles, the volume, and the temperature a distribution $p_{\text{can}}(\cdot|N,V,T)$ over the set of microstates from which the decoded state is drawn, $h^{-1}_{\text{dec}}\sim p_{\text{can}}(\cdot|N,V,T)$.
The above objective depends on the choice of distance $d$, which incorporates the geometry of the state space and any desired biases about what is “important” in the system; it may differ from system to system. In what follows, we measure accuracy in terms of the mutual information $I$ shared between the reduced model prediction and the full state. This abstracts away from the geometry of the space and characterizes the similarity of variables in terms of how predictive they are of each other, rather than how close they are, in a system-independent way. This has an important consequence: as a result of the data processing inequality, we can show that our objective function simplifies to one depending only on $h_{\text{enc}}$:
$\min_{h_{\text{enc}}}$ $\quad\dim h_{\text{enc}}(X_{t})-\beta I\big(h_{\text{enc}}(X_{t}),X_{t+\Delta t}\big)$ (2)
The steps to this result are shown in detail in the Methods. This simplification is dramatic: instead of simultaneously learning the encoding function, evolution operator, and decoder, the only step we need to care about is the encoding. Consequently, we view model reduction in the rest of this paper as a task of finding the optimal reduced variables with which to represent our system. The evolution rule $\mathcal{S}$ and decoder $h^{-1}_{\text{dec}}$ are induced by the choice of $h_{\text{enc}}$, and we may view their identification as a secondary, downstream task to be performed after the proper variables have been found.
Before we proceed, we still need to specify the dimensionality reduction. This is subtle, as in different contexts different measures may seem natural, for instance, the entropy for discrete variables, or the number of components for continuous variables. We achieve our dimensionally-reduced variables with a two-step construction. First, we use a term explicitly forcing the reduced representation to lose information about the full state, $\min I(h(X_{t}),X_{t})$ (for notational simplicity, we refer to $h_{\text{enc}}$ as $h$). The minimal objective found in this way
$\min_{h}\mathcal{L}_{\text{IB}}=I(X_{t},h(X_{t}))-\beta I(X_{t+\Delta t},h(X_{t})),$ (3)
is known as the information bottleneck (IB) [38, 39]. It (and related variants [42, 43]) has been used in a variety of contexts, from document clustering [42] to neural codes [44] to (most relevant for us) dynamical systems [36, 45, 46], as well as in deep-learning contexts [47, 41, 48].
@@
The IB objective ((3)) is a functional over stochastic functions, which may be represented by a conditional probability distribution $h(x_{t})\sim p(h|x_{t})$ (we denote random variables by uppercase $X_{t}$ and their values by lowercase $x_{t}$). Crucially, the parameter $\beta$ controls the trade-off between compression and prediction. For small $\beta$ the compression term dominates and the optimal encoder is trivial, losing all information about the system. For intermediate $\beta$ the compression term does not allow $X_{t}$ to be completely represented by $H_{t}$, so features of $X_{t}$ must compete to pass through to the encoding variable (Fig. 1b). These selected features are reflected in the IB encoder
@@
which provides the optimal trade-off between retained information and predictability [37]. In the optimal encoder, information about $X_{t}$ has been removed both by reducing the dimensionality, as well as by injecting noise into $h_{t}$. Increasing $\beta$ can impact both of these, either by increasing the dimensionality to learn new features, or by reducing noise to better identify the features that have already been found. In order to “perfectly” identify the features given a fixed set of components, in the second step we modify the IB encoder by discarding the part attributable to noise alone. This effectively increases the coupling of the reduced variables to the already-learned features. We shall see that this procedure isolates deterministic, interpretable reduced variables in both discrete and continuous settings.
Figure 1: Interpretable dynamical variables in reduced models via the information bottleneck. (a) In general model reduction aims to both learn an encoding and reduced dynamics, so that the reduced variable serves as a useful proxy that may be decoded for an estimate of the full system. Requiring the reduced variable to only be informative of the full state decouples these two steps (see SI Section 2). (b) The information bottleneck compresses high-dimensional state variables $x_{t}$, into simpler encoding variables $h_{t}$ with a controllable trade-off between the degree of compression and the predictive power about the system’s future. With deep neural networks, the encoding can be computed directly from data of observed fluid flows (left) or biological datasets, such as fluorescently labeled bacteria colonies (right). In general, the state of the variable $x_{t}$ may comprise time-lagged variables of the intensity field, $x_{t}=\{I_{t},I_{t+\Delta t}\}$ (right). The amount of compression is determined by the “width” of the bottleneck $\beta$ (see (3)). The resulting compressed, or encoded, variables $h_{t}$ represent collective variables most predictive of the system’s future. (c) Schematic evolution of the encoder $p(h_{t}|x_{t})$ for varying compression strength $\beta$. For low $\beta$ (high compression), the encoder is trivial and forgets everything about the input $x_{t}$. After the first IB transition at $\beta_{1}$, the encoder becomes non-trivial by gaining some dependence on $x_{t}$; some features of the input are able to pass through the bottleneck. At subsequent IB transitions, additional features are learned. (d) The point spectrum of the transfer operator contains several slowly decaying modes (red). We show that the most predictive variables that IB systematically extracts correspond to the slowest eigenfunctions of the transfer operator, associated to eigenvalues $\Lambda_{i}$ with $|\Lambda_{i}|\approx 1$. In fluid flows, the slowest-decaying eigenfunctions typically represent large-scale coherent patterns of the flow field, while faster-decaying eigenfunctions correspond to variations over shorter length scales. In general the system may exhibit an essential spectrum, however for noisy systems this is small (see main text).
II The optimal encoder in terms of the transfer operator
@@
In any realistic experimental setting, the presence of noise or uncertainty means we cannot predict precisely the future state of a system but instead only a likely distribution of future states. Our prediction of the state at $\Delta t$ in the future is then represented mathematically as $p(x_{t+\Delta t}|x_{t})$, the probability of observing state $x_{t+\Delta t}$ given the current state $x_{t}$. For Markovian, or “memoryless” dynamics, this conditional probability distribution completely characterizes the dynamics of the system, and determines how probability distributions evolve in time. In this work we assume that the systems we consider are (or can be made) Markovian; for a detailed discussion of how to measure the memory and how it changes our below results, see SI Sections 3 and 4D. The evolution of probability distributions can then be understood as the action of a linear transfer operator $U^{\Delta t}$ on probability distributions:
$p_{X_{t+\Delta t}}(x_{t+\Delta t})=\int p(x_{t+\Delta t}|x_{t})p_{X_{t}}(x_{t})\text{d}x_{t}\equiv U^{\Delta t}[p_{X_{t}}]$ (5)
$U^{\Delta t}$ can be decomposed as follows (using standard notation from quantum mechanics and explained in the SI):
[…]
@@
Our key observation is that the optimal IB encoder in (4), which we will use as a “filter” to extract relevant features, can be expressed in terms of the eigenvalues $\lambda_{n}$ and left eigenfunctions $\phi_{n}$ of (the generator of) $U$,
@@
where $f_{n}(h_{t})$ are factors that do not depend on $x_{t}$. For an outline of the mathematical steps leading to this see Methods, as well as SI Section 4. The combined effect of the factors in the exponent determines what features the encoder learns about the state $x_{t}$. In general, there may be a large number of non-zero factors $f_{n}$ so that the learned features are difficult to interpret in terms of individual eigenfunctions $\phi_{n}$. However, things become simple in the limit of small $\beta$, or high compression. When $\beta$ is small the encoder is trivial: $p(h_{t}|x_{t})=p(h_{t})$ so the value $h_{t}$ is assigned at random with no regard to the state of the system. No feature has been learned, and all factors $f_{n}$ are equal to zero. As $\beta$ is increased, the encoder undergoes a series of transitions at $\beta=\beta_{1}<\beta_{2}<\beta_{3}$... where new features are allowed to pass through the bottleneck (Fig. 1c) [51, 52, 53, 54]. The first transition happens at a finite value of $\beta_{1}$ when the first, most predictive, feature is learned.
@@
This is our main mathematical result, which we derive by considering a perturbative expansion of the IB objective for small $f_{n}$. A proof of Eq. 8 with clearly specified technical assumptions may be found in Section 4 of the SI. One notable assumption underpinning this result, in addition to Markovianity, is that the system is in a statistical steady state, so that $x_{t}$ is sampled from the steady state distribution. When this assumption is violated, eigenfunctions of the transfer operator may cease to carry information about the system’s dynamics and instead its pseudospectrum may become relevant [55]; see SI Section 4D.
@@
The above result shows that in the limit of high compression the encoder’s dependence on $x_{t}$ is given by the first left eigenfunction $\phi_{1}(x_{t})$, which is the slowest-varying function of the state under dynamics given by $U$. Therefore, Eq. 8 makes precise the intuitive statement that slow features are the most relevant for predicting the future. Our analytical result, while applying only to the dominant eigenfunction, is valid for arbitrary, including non-Gaussian, variables.
@@
We further observe numerically that this picture holds true more generally: at successive IB transitions, the learned features correspond to successive modes of the transfer operator. This is consistent with the exact results known for the specific case of Gaussian IB, where the encoder learns successive eigenvectors of a matrix related to the covariance of the joint $X_{t},X_{t+\Delta t}$ distribution at each IB transition [53].
Due to the smallness of $f_{1}$ at the transition, the optimal IB encoder couples only weakly to the eigenfunction $\phi_{1}$, so that the dynamics of $h_{t}$ will be dominated by noise. Given the maximal number of available bits in memory, or variables used in an analytical model, the IB solution will thus not make the best use of the resources (IB is optimal with respect to the amount of information retained, but not the size of its representation [43]). However, we can use the optimal IB encoding as a filter to identify $\phi_{1}$, and then “distill” it by discarding the noisy part of the encoding, effectively increasing the coupling between $h_{t}$ and $\phi_{1}(x_{t})$ while keeping the number of components fixed. In practice, this is done by taking the most likely value for the reduced state, $h_{t}=\arg\max_{h}p(h|x_{t})$ (see SI Section 4E). Similar procedures are performed in other probabilistic dimensionality reduction approaches, e.g. probabilistic principal component analysis [56]. There one finds a stochastic mapping from the full state $x$ to the reduced representation $h$, and then discards the noise to obtain the deterministic reduction.
Together this lays the foundation of an operational prescription for model reduction which is built on IB for identifying relevant, slow, features. As we show later, this insight can be leveraged to systematically learn these slow variables directly from data with neural networks [41] which in turn may be used as inputs to an equation learning pipeline.
III Information decay and the spectrum of the transfer operator